### v1

Abstract
Current studies have exposed the risk of Large Language Models (LLMs) generating harmful content by jailbreak attacks. However, they overlook that the direct generation of harmful content from scratch is more difficult than inducing LLM to calibrate benign content into harmful forms. In our study, we introduce a novel attack framework that exploits AdVersArial meTAphoR (AVATAR) to induce the LLM to calibrate malicious metaphors for jailbreaking. Specifically, to answer harmful queries, AVATAR adaptively identifies a set of benign but logically related metaphors as the initial seed. Then, driven by these metaphors, the target LLM is induced to reason and calibrate about the metaphorical content, thus jailbroken by either directly outputting harmful responses or calibrating residuals between metaphorical and professional harmful content. Experimental results demonstrate that AVATAR can effectively and transferable jailbreak LLMs and achieve a state-of-the-art attack success rate across multiple advanced LLMs.
[…]
from Benign import Toxic: Jailbreaking the Language Model via Adversarial Metaphors
Warning: This paper contains potentially harmful content.
Yu Yan1,2, Sheng Sun1, Zenghao Duan1,2, Teli Liu3, Min Liu5, Zhiyi Yin1, Qi Li4, Jingyu Lei4, 1Institute of Computing Technology, Chinese Academy of Sciences, Beijing, China 2University of Chinese Academy of Sciences, Beijing, China 3People’s Public Security University of China, Beijing, China 4Tsinghua University, Beijing, China 5Zhongguancun Laboratory, Beijing, China {yanyu24z, sunsheng, duanzenghao24s, liumin}@ict.ac.cn, {qli01, leijy}@tsinghua.edu.cn
1 Introduction
(a) Generating harmful content from scratch
(b) Calibrating the metaphor to harmful content
Figure 1: Illustration of inducing target LLM (GPT-4o) for harmful metaphor analysis (Cook the dish $\rightarrow$ Build a bomb ) and using the target LLM to calibrate the metaphorical content for jailbreaking.
Large Language Models (LLMs) Yi et al. (2024); Achiam et al. (2023) have become increasingly prevalent across various domains, such as content generation Wang et al. (2024a), programming Grattafiori et al. (2023), and professional question answering Thirunavukarasu et al. (2023).
However, recent studies Zou et al. (2023); Yi et al. (2024); Zhu et al. (2023); Chang et al. (2024); Zeng et al. (2024) have revealed attacks targeting LLM safety alignment mechanisms, known as jailbreak attacks, which aim to break or bypass the LLM’s built-in content filtering and safety protections, generating biased and harmful content. Among them, black-box jailbreak attacks Yi et al. (2024) present a more widespread threat, as these attacks can be carried out without access to the internal LLMs and exhibit transferability across different LLMs. Existing attacking techniques mainly focus on language rewriting Chao et al. (2023); Yong et al. (2023); Mehrotra et al. (2023); Ding et al. (2024); Yuan et al. (2024); Handa et al. (2024) or predefined templates Yao et al. (2024); Yu et al. (2023); Wei et al. (2023); Jiang et al. (2024); Li et al. (2023a) for effective adversarial attacks. However, these end-to-end attack paradigms, which generate harmful content from scratch, often fail to decouple malicious intent from the harmful query, making them more detectable and increasingly ineffective as LLMs evolve. This could be interpreted as two critical aspects being neglected: 1) the effectiveness of weaponizing the benign auxiliary data for manipulating LLM’s goal prioritization and 2) the risk of rewriting the content into harmful forms by accumulating residual harmful knowledge.
To advance the understanding of black-box jailbreak attacks, we innovatively explore the novel paradigm that progressively transforms benign metaphors into harmful outputs, rather than generating them from scratch. As these benign metaphors share fundamental similarities with harmful concepts in logical reasoning chains, causal relations, or functional mechanisms, when LLMs analyze these metaphors, the models’ inherent cognitive mapping process inadvertently reveals knowledge applicable to the harmful target domain. As shown in Figure 1, to generate responses for the harmful query "Build a Bomb", we introduce the benign metaphor "Cook the Dish", since both require precise control over execution steps and careful composition of ingredients. Subsequently, once we have accumulated enough information in the interaction, we can calibrate these metaphorical analyses into the professional response toward the harmful task, i.e., "Build a Bomb" for jailbreaking.
Inspired by this, we propose a novel attack framework for black-box harmful content generation using AdVersArial meTAphoR (AVATAR). Specifically, we introduce Adversarial Entity Mapping (AEM) to identify and select appropriate metaphors based on the imagination of crowdsourced LLMs. Then, we design Metaphor-Driven Thinking (MDT) to drive the first-stage jailbreak attempt, where the target LLM analyzes harmful tasks through metaphorical reasoning. Finally, the target LLM is induced to calibrate the residuals between metaphorical and professional answers to the given query for the second-stage jailbreak attempt. We conduct extensive experiments to demonstrate the effectiveness and transferability of AVATAR. Our major contributions [^1] Our code is available at:
https://anonymous.4open.science
are:
  • We propose a novel perspective that calibrates benign content into harmful for jailbreaking, instead of generating from scratch. Based on this, we present AdVersArial meTAphoR (AVATAR), which leverages metaphors to drive LLM for harmful knowledge output while maintaining intentions innocuous.
  • We introduce Adversarial Entity Mapping (AEM), which generates metaphors using the crowdsourced LLMs and conducts Metaphor-Driven Thinking (MDT) to induce target LLMs to generate harmful content.
  • We analyze the mechanisms behind metaphor attacks and explore potential defense methods. Extensive experiments demonstrate that AVATAR achieves state-of-the-art ASR (over 92% on GPT-4o within 3 retries), and successfully utilizing powerful LLMs such as ChatGPT-o1 for harmful content generating.
Figure 2: Overview of our AVATAR, which is the black-box attack method without training, involving two main steps: First, Adversarial Entity Mapping identifies the appropriate metaphors by balancing the effectiveness of harmful content with toxicity concealment by crowdsourced models. Then, Metaphor-Driven Thinking nests metaphors into interactions and induces the target model to generate harmful output from the metaphorical analysis.
2 Problem Statement
In this study, we formalize our jailbreak attack in AVATAR as a metaphorical attack, which can be simplified as a directed semantic graph construction problem. Specifically, given the set of entities $\mathcal{E}=\mathcal{E}_{h}\cup\mathcal{E}_{b}$, where $e\in\mathcal{E}$ is the related entity in either harmful or benign task completion process, ${\mathcal{E}_{h}}$, ${\mathcal{E}_{b}}$ are the set of harmful entities and benign entities, $\mathcal{R}$ is the set of semantic relations, where $r\in\mathcal{R}$ captures the semantic relations between entities in $\mathcal{E}$. The goal of the metaphorical attack is to jailbreak the target LLM $\mathcal{G}_{\text{target}}$ to extend the harmful entities ${\mathcal{E}_{h}}$, and steal the set of relations ${R_{h}}=\{(u,v)\mid u,v\in\mathcal{E}_{h}\}$, which captures semantic relations among harmful entities $\mathcal{E}_{h}$ by using $\mathcal{E}$, ${\mathcal{R}_{b}}=\{(u,v)\mid u,v\in\mathcal{E}_{b}\}$ and ${\mathcal{R}_{m}}=\{(u,v)\mid u\in\mathcal{E}_{b},v\in\mathcal{E}_{h}\}$.
3 Methodology
In this section, we introduce AdVersArial meTAphoR (AVATAR) as shown in Figure 2, which first generates the benign content, and then calibrates the residuals between benign and harmful answers to the given query for jailbreaking. The role settings are in Appendix.
3.1 Adversarial Entity Mapping
To adaptively identify metaphors, we propose the Adversarial Entity Mapping (AEM) approach, which can stably discover appropriate entities as the initial seed for harmful content generated by model crowdsourcing, as shown in Figure 3.
Toxic Entity Extraction.
Generally, harmful tasks imply some necessary prerequisites and can be represented as some entities. We extract them from the harmful task to enable the subsequent deep logical metaphor matching. The attacker model $\mathcal{G}_{\text{attacker}}$ is used to extract the original toxic entity $E_{\text{ori}}$ and its associated sub-entities $\mathcal{E}_{\text{ori}}$ from the harmful query $Q_{\text{ori}}$. Formally, given $Q_{\text{ori}}$, the original toxic entity $E_{\text{ori}}$ is extracted and concatenate with the template $P_{\text{S}}$ to extract harmful sub-entities $\mathcal{E}_{\text{ori}}$ using $\mathcal{G}_{\text{attacker}}$:
$\mathcal{E}_{\text{ori}}=\mathcal{G}_{\text{attacker}}(P_{\text{S}}\oplus E_{\text{ori}}),$ (1)
where $\mathcal{E}_{\text{ori}}=\{e_{\text{ori}}^{0},e_{\text{ori}}^{1},...\}$ is the original key sub-entity set for $E_{\text{ori}}$. To further strengthen the logical association between sub-entities and the original harmful entity, the attacker model $\mathcal{G}_{\text{attacker}}$ is then utilized to select $k$ entities to refine $\mathcal{E}_{\text{ori}}$ as $\{e_{\text{ori}}^{0},e_{\text{ori}}^{1},...e_{\text{ori}}^{k}\}$.
Metaphor Entity Identifying.
After describing the harmful target through a set of core harmful entities, we identify corresponding parallel entities to metaphorically represent them. To achieve this, we introduce high-temperature and model crowdsourcing strategy to discover mapping entities.
The high-temperature setting increases the variability and creativity of the outputs, allowing us to explore more metaphors. Meanwhile, the model crowdsourcing strategy ensures that more suitable metaphors are steadily obtained from various knowledge perspectives before jailbreaking the target model. Formally, given the original entity $E_{\text{ori}}$ and its associated sub-entities $\mathcal{E}_{\text{ori}}$, the $i$-th selected tool model $\mathcal{G}_{\text{tool}}^{i}$ in the model crowdsourcing pool is used to imagine mapping entity $E_{\text{map}}^{i}$ and its sub-entities $\mathcal{E}_{\text{map}}^{i}$ based on cause $C_{\text{map}}^{i}$ (a text that explains the cause for the mapping):
$(E_{\text{map}}^{i},\mathcal{E}_{\text{map}}^{i},C_{\text{map}}^{i})$ $=\mathcal{G}_{\text{tool}}^{i}(E_{\text{ori}}\oplus\mathcal{E}_{\text{ori}}\oplus P_{\text{M}},\sigma),$ (2)
$\mathcal{S}_{\text{map}}$ =$=\{(E_{\text{map}}^{i},\mathcal{E}_{\text{map}}^{i},C_{\text{map}}^{i})\}_{i=1}^{M},$
where $\sigma$ is the temperature parameter and $P_{\text{M}}$ is the task prompt for entity mapping. $\mathcal{S}_{\text{map}}$ is the set consisting of triples from $M$ tool models.
Minimum Toxicity Metaphor.
After generating a set of candidate mapping entities, our final step for AEM is to select the minimal toxicity entities that reflect the internal relations of the original harmful entities, while avoiding the exposure of harmful intent. To achieve this balance, we introduce Internal Consistency Similarity (ICS) and Conceptual Disparity (CD) to filter the candidate set $\mathcal{S}_{\text{map}}$, aiming to maximize the toxicity of harmful content while minimizing the risk of triggering LLM’s safety alignment mechanisms.
Internal Consistency Similarity (ICS). To ensure that the mapping entities $E_{\text{map}}$ retain coherent and meaningful relations with the original entities $E_{\text{ori}}$, we measure the similarity of their internal entity relations. ICS is calculated using two unified sets, the original entity with its sub-entities $\mathcal{U_{\text{ori}}}=\{E_{\text{ori}}\}\cup\mathcal{E}_{\text{ori}}$ and mapping entity with its sub-entities $\mathcal{U_{\text{map}}}=\{E_{\text{map}}\}\cup\mathcal{E}_{\text{map}}$, defined as:
ICS $=\text{sim}(\mathbf{M}_{\text{ori}},\mathbf{M}_{\text{map}}),$ (3)
$\mathbf{M}_{\text{ori}}$ $=\left[\text{sim}(\mathbf{v}_{a},\mathbf{v}_{b})\right]_{a,b\in\mathcal{U}_{\text{ori}}},$
$\mathbf{M}_{\text{map}}$ $=\left[\text{sim}(\mathbf{v}_{a},\mathbf{v}_{b})\right]_{a,b\in\mathcal{U}_{\text{map}}},$
where $\mathbf{M}_{\text{ori}}$ and $\mathbf{M}_{\text{map}}$ are internal entity similarity matrices, $\mathbf{v}\in\mathbb{R}^{h}$ is the semantic representation of entity calculated by the embedding model, e.g., BGE-M3 Chen et al. (2024), with $h$ hidden state dimension, $\text{sim}(\mathbf{v},\mathbf{u})$ is cosine similarity.
Figure 3: Illustration of Adversarial Entity Mapping, which creates adversarial metaphors via crowdsourcing.
Conceptual Disparity (CD). To further ensure that the mapping entity $E_{\text{map}}$ remains distinct from harmful entity $E_{\text{ori}}$ while retaining enough relevance for generating valid responses, we measure their external entity disparity. CD averages the similarity between the sets $\mathcal{U_{\text{ori}}}$ and $\mathcal{U_{\text{map}}}$ as:
CD $=\frac{\sum\mathbf{M}_{\text{cross}}}{|\mathbf{M}_{\text{cross}}|},$ (4)
$\mathbf{M}_{\text{cross}}$ $=\left[\text{sim}(\mathbf{v}_{a},\mathbf{v}_{b})\right]_{a\in\mathcal{U}_{\text{ori}},b\in\mathcal{U}_{\text{map}}},$
where $\mathbf{M}_{\text{cross}}$ is entity cross similarity matrix of $\mathcal{U_{\text{ori}}}$ and $\mathcal{U_{\text{map}}}$, which captures similarities between corresponding elements.
Optimization Goal. In Minimum Toxicity Metaphor, we find appropriate mapping entities $E_{\text{map}}$ and $\mathcal{E}_{\text{map}}$ for entity-level jailbreak by balancing the effectiveness of harmful content (ICS) with toxicity concealment (CD). Therefore, we introduce a sigmoid transformation, which attains higher values when ICS and CD are well-balanced, we formalize the optimized goal of MTM as:
[…]
3.2 Metaphor-Driven Thinking
We introduce benign metaphors as cognitive payloads to exploit the similarities and differences between harmful targets and their metaphorical counterparts. This approach enables the systematic extraction of task-completion knowledge from target LLMs through analogical reasoning.
Metaphorical Payload Setup.
We adaptively load the adversarial metaphors into a series of interaction queries for initialization, consisting of based queries and adaptive queries.
Based queries consist of two fixed queries, $Q_{\text{ctx}}$ presents the context of metaphors to the LLM, and $Q_{\text{det}}$ induces the LLM for detailed explaining metaphors, defined as:
$Q_{\text{ctx}},Q_{\text{det}}$ $=P_{\text{ctx}}\oplus E_{\text{ori}}\oplus E_{\text{map}}\oplus\mathcal{E}_{\text{ori}}\oplus\mathcal{E}_{\text{map}}$ (6)
$\quad\oplus C_{\text{map}}\oplus\mathbf{M}_{\text{ori}}\oplus\mathbf{M}_{\text{map}}\oplus\mathbf{M}_{\text{cross}},P_{\text{det}},$
where $P_{\text{ctx}}$, $P_{\text{det}}$ are templates, the metaphor context ($E_{\text{ori}}$, $E_{\text{map}}$, $\mathcal{E}_{\text{ori}}$, $\mathcal{E}_{\text{map}}$, $C_{\text{map}}$, $\mathbf{M}_{\text{ori}}$, $\mathbf{M}_{\text{map}}$, $\mathbf{M}_{\text{cross}}$) is attached with $P_{\text{ctx}}$ in JSON format.
Adaptive queries $\mathcal{Q}_{\text{ext}}$ are generated by the attacker model to further ensure the success of jailbreak according to the given attack target and corresponding metaphors, defined as:
$\mathcal{Q}_{\text{ext}}$ $=\mathcal{G}_{\text{attacker}}(P_{\text{ext}}\oplus E_{\text{ori}}\oplus E_{\text{map}}\oplus$ (7)
$\quad\mathcal{E}_{\text{ori}}\oplus\mathcal{E}_{\text{map}}\oplus C_{\text{map}}\oplus N),$
where $P_{\text{ext}}$ is the task prompt for jailbreak query extending, $N$ is the total query for generating.
Initial interaction queries $\mathcal{Q}_{\text{init}}$ combine based queries and adaptive queries. Specifically, to further ensure the effectiveness of adaptive queries, $\mathcal{Q}_{\text{ext}}$ is filtered as $\mathcal{Q}_{\text{ext}}^{*}$ with top-$k$ queries most relevant to original harmful query $Q_{\text{ori}}$. By integrating $Q_{\text{ctx}}$, $Q_{\text{det}}$, and $\mathcal{Q}_{\text{ext}}^{*}$, we obtain initial interaction queries $\mathcal{Q}_{\text{init}}$, defined as:
$\mathcal{Q}_{\text{init}}=\{Q_{\text{ctx}},Q_{\text{det}}\}\cup\text{SortDesc}(\mathcal{Q}_{\text{ext}}^{*},\text{sim}(\cdot,Q_{\text{ori}})),$ (8)
$\mathcal{Q}_{\text{ext}}^{*}=\underset{\mathcal{Q^{\prime}}\subseteq\mathcal{Q}_{\text{ext}},|\mathcal{Q^{\prime}}|=k}{\arg\max}\left\{\mathbb{E}\left[\text{sim}\left(\mathcal{Q^{\prime}},Q_{\text{ori}}\right)\right]\right\},$ (9)
where $\text{SortDesc}(\cdot,\cdot)$ sorts the queries in descending order of their toxic-aware similarity. $\text{sim}(\mathcal{Q},Q_{\text{ori}})$ is query similarity calculation based on toxic-aware embedding tool. $\mathcal{Q^{\prime}}$ represents any subset of $\mathcal{Q}_{\text{ext}}$ that includes exactly $k$ queries, and $\mathbb{E}[\cdot]$ is the expected value of the similarity scores between the queries in $\mathcal{Q^{\prime}}$ and $Q_{\text{ori}}$.
Adversarial Human-like Interaction Optimization.
To ensure the success of jailbreak, we incorporate human social influence strategies Zeng et al. (2024); Wang et al. (2024b) to refine queries based on feedback from LLM’s response. Besides, another critical consideration is the management of the interaction state, i.e., the conversation history.
As suggested in previous work Yang et al. (2024), the relevance of LLM’s responses to the original harmful target should gradually increase during the interaction. Formally, consider a query $Q_{t}$ in the $t$-th round interaction, the response from the target model $\mathcal{G}_{\text{target}}$ with/without the historical context up to the $t$-th round is defined as $R_{t}$, $R_{t}^{\prime}$:
$R_{t},R_{t}^{\prime}=\mathcal{G}_{\text{target}}(Q_{t}\mid\mathcal{C}),\mathcal{G}_{\text{target}}(Q_{t}),$ (10)
where $\mathcal{C}=\{(Q_{j},R_{j})\mid 1\leq j\leq t-1\}$ is the conversation history. $R_{t}^{\prime}$ serves as a baseline to evaluate the effectiveness of historical context for harmful response. By comparing the $R_{t}$ and $R_{t}^{\prime}$, the management of the interaction state is determined by the following $Q$-$R$ similarity conditions:
  • If $\text{sim}(R_{t},Q_{\text{ori}})$ $>$ $\text{sim}(R_{t-1},Q_{\text{ori}})$ and $\text{sim}(R_{t},Q_{\text{ori}})>\text{sim}(R_{t}^{\prime},Q_{\text{ori}})$, then append the response $R_{t}$ to $\mathcal{C}$ for $t+1$ round.
  • If $\text{sim}(R_{t},Q_{\text{ori}})\leq\text{sim}(R_{t-1},Q_{\text{ori}})$ and $\text{sim}(R_{t}^{\prime},Q_{\text{ori}})\geq\text{sim}(R_{t-1},Q_{\text{ori}})$ , then turn back to $t-1$ round historical interaction.
  • Otherwise, refine $Q_{t}$ with human social influence strategies for the next step re-evaluation.
To refine queries, we identify the harmful type of the given queries and select a human social influence strategy randomly from the top-5 most effective strategies Zeng et al. (2024) to assist in rewriting. Within max_round iteration for interaction adjusting, we seek to obtain the harmful response. To further induce target LLM, we manipulate it to calibrate the response in each round into the professional and harmful response according to the metaphorical relations. This metaphor calibration is an effective residual mechanism to fully leverage the existing context for jailbreaking.
Figure 4: Illustration of Metaphor-Driven Thinking, which loads adversarial metaphors into a series of queries and adaptively adjusts queries according to LLMs’ feedback.
4 Experiments
In this section, we assess the feasibility and effectiveness of metaphorical attacks by conducting experiments on several widely used LLMs.
4.1 Experiment Settings
Datasets and Models.
We select the standard and contextual behavior with a total of 240 samples from HarmBench Mazeika et al. (2024) and the top-50 most toxic samples Yang et al. (2024) from AdvBench Zou et al. (2023) to evaluate the attack ability of our AVATAR framework, both of which are widely used in the fields Ding et al. (2024); Zeng et al. (2024); Yu et al. (2023). As for the target model, we select the advanced LLMs, Qwen2.5-7B, Llama3-8B, GPT-4o-mini, and GPT-4o, which are commonly used for current AI applications.
Evaluation Metrics.
We employ two Attack Success Rate metrics for measuring attack effectiveness against different harmful queries: ASR-KW Zou et al. (2023) and ASR-GPT Mazeika et al. (2024); Yang et al. (2024); Li et al. (2024). Specifically, ASR-KW (Keyword-based Attack Success Rate) mainly evaluates the methods’ bypassing ability while lacking consideration of the toxicity. To mitigate the issue of false positives in keyword detection approaches and address the limitations of toxicity classifiers Yu et al. (2023); Hartvigsen et al. (2022), both of which overlook the task relevance of the generated content. We utilize the GPT-4o and predefined criteria prompt from Harmbench Mazeika et al. (2024) to evaluate the relevance and potential harmfulness of model outputs as the ASR-GPT (GPT-based Attack Success Rate).
Baseline Methods.
We use the following 6 baselines for comparison: AutoDAN Zhu et al. (2023), PAIR Chao et al. (2023), TAP Mehrotra et al. (2023), CoA Yang et al. (2024), SelfCiper Yuan et al. (2024), DrAttack Li et al. (2024).
Method Qwen2.5-7b-Instruct Llama3-8b-Instruct GPT-4o-mini GPT-4o
ASR-KW ASR-GPT ASR-KW ASR-GPT ASR-KW ASR-GPT ASR-KW ASR-GPT
AutoDAN 92.92 63.33 87.92 77.08 - - - -
PAIR 64.58 43.75 40.42 11.25 50.83 42.92 41.25 35.83
TAP 77.50 66.25 51.25 25.42 68.75 52.92 62.08 47.08
COA 83.75 65.83 58.33 32.50 75.83 63.75 65.00 50.42
SelfCipher 100.00 54.58 100.00 57.92 100.00 77.92 100.00 63.33
DrAttack 95.42 83.33 77.08 42.50 94.58 81.67 82.08 76.67
AVATAR (ours) 100.00 100.00 100.00 97.08 100.00 95.83 97.50 92.08
Table 1: Experimental ASR-KW (%) and ASR-GPT (%) of various methods across four mainstream LLMs on Harmbench. The best results are highlighted in bold. The second best results are highlighted in underline.
Figure 5: Transfer attack performance (ASR-GPT, %) of AVATAR on Harmbench. The attack is conducted by using adversarial prompts whose effectiveness is verified on affordable LLMs (Qwen2-7B, Llama3-8B, and GPT-4o-mini). Fixed template means we only load the adversarial metaphor on based queries ($Q_{\text{ctx}}$, $Q_{\text{det}}$) for induction. Adaptive opt. means we introduce adaptive queries and adversarial human-like interaction optimization.
Detailed experimental settings are in Appendix.
4.2 Experiment Results
Baseline Comparison.
We perform the attack methods in 3 retries to jailbreak LLMs and compare their performance with our AVATAR. As shown in Table 1, AVATAR demonstrates superior performance across four mainstream LLMs. Among them, AVATAR achieves 100% ASR-GPT on Qwen2.5-7B and Llama3-8B, outperforming the second-best performance by 16.67% and 22.92% respectively. Furthermore, AVATAR can effectively jailbreak closed-source models GPT-4o and GPT-4o-mini, achieving 92.08% and 95.83% on ASR-GPT, surpassing the second-best performance by 15.41% and 14.16% respectively. A further case study is in Appendix demonstrates that AVATAR can further jailbreak powerful LLMs such as ChatGPT-o1 and Claude-3.5.
Compared to current methods generating harmful content from scratch, AVATAR achieves higher ASR by utilizing benign metaphors as camouflage. The following is further analyzed in detail: 1) Building upon multi-turn interaction jailbreak methods (PAIR, TAP, COA) that merely adjust expression techniques, AVATAR can use metaphors to induce models to generate harmful content more willingly. 2) In contrast to methods like DrAttack and SelfCipher that jailbreak the models by distracting model attention, e.g., sub-prompt splitting and text encrypting, AVATAR introduces harmless auxiliary data to transform attack tasks into metaphorical interpretation tasks, thus naturally inducing LLMs to generate harmful content.
Transfer Attack.
To evaluate the transferability of AVATAR, we conduct transfer attack experiments on Harmbench across 10 mainstream LLMs. We obtain the attack prompts in the metaphor-driven thinking stage from three affordable LLMs for experiments: 1) Llama3-8B, 2) Llama3-8B and Qwen2.5-7B, 3) Llama3-8B, Qwen2.5-7B, and GPT-4o-mini. The results are shown in Figure 5, demonstrating the transferability of our AVATAR.
With attacking by fixed template from Llama3-8B, AVATAR can achieve notable ASR-GPT over 75.00% in Gemma2-9B, GLM3-6B, and Mistral-7B. When further introducing the adaptively optimized prompt, the performance significantly improves, e.g., Qwen2.5-32B (10.90$\uparrow$), Qwen2-72B (24.20$\uparrow$), Llama3.1-8B (30.40$\uparrow$). Furthermore, simultaneously leveraging adversarial prompts from three models for transfer attacks shows superior performance achieving ASR-GPT above 90.00% on most target models (e.g., 98.33% on GLM3-6B, 99.17% on Mistral-7B). This demonstrates that AVATAR can generate highly transferable adversarial prompts, highlighting the effectiveness of the adversarial metaphor.
Ablation Study.
To verify the effectiveness of each component in AVATAR, we conduct the ablation study as shown in Table 2. We further analyze the influence of different metaphor selection strategies in Appendix.
To validate our adversarial metaphor method, we try to jailbreak LLMs by removing Adversarial Entity Mapping (w/o Adv.Map), only using Adaptive Queries for attack initialization and Adversarial Human-like Interaction Optimizaion for prompt adjusting. The significant drop in ASR-GPT (39.04% in GPT-4o-mini, 39.51% in GPT-4o) is observed without Adv. Map, which underscores the effectiveness of introducing benign metaphors for harmful content generating while concealing harmful intent.
Variants ASR-GPT (%)
Qwen2.5-7B GPT-4o-mini GPT-4o
AVATAR 100.00 95.83 92.08
[…]
w/o M.Crowd 83.28 16.72↓ 82.63 13.20↓ 74.40 17.68↓
w/o HI.Opt. 91.72 8.28↓ 89.27 6.56↓ 87.67 4.41↓
Table 2: Experimental ASR-GPT (%) of Ablation Study for AVATAR variants on Harmbench, averaging results from 3 repeated experiments.
Figure 6: Iteration statistics for successful jailbreak queries across four mainstream LLMs on Harmbench.
To validate the usefulness of crowdsourced models’ metaphor identifying, we remove Model Crowdsourcing and only use the attacker model for Adversarial Entity Mapping (w/o M.Crowd). The noticeable decrease in ASR (13.20% in GPT-4o-mini, 17.68% in GPT-4o) indicates that one single LLM’s imagination is not robust enough to identify appropriate metaphors for jailbreaking. On one hand, the model crowdsourcing strategy provides diverse mapping options, on the other hand, it avoids attack failures caused by a single LLM’s rejection of metaphor identification.
To validate our prompt optimization method, we remove Adversarial Human-like Interaction Optimization (w/o HI.Opt.), and only use the based queries ($P_{\text{ctx}}$, $P_{\text{det}}$). While the impact on ASR is less pronounced (6.56% in GPT-4o-mini, 8.28% in Qwen-2.5), it is crucial for AVATAR’s stealth. HI.Opt. maintains performance by nesting harmful intents within more natural interactions, allowing the attack to remain ostensibly innocuous even when metaphor choices are imperfect.
Efficiency of Jailbreak.
To verify the efficiency of AVATAR, we compare its iteration counts for successful jailbreaks with those of CoA and AVATAR’s MDT module on Harmbench. As shown in Figure 6, AVATAR requires significantly fewer iterations (4.9 on average) than CoA (18.3 on average), highlighting AVATAR’s superior efficiency improvement from metaphorical induction. Furthermore, we can observe that the average iteration for jailbreaking is significantly reduced (7.6$\rightarrow$4.9) by introducing metaphor calibration. The metaphor calibration module can potentially induce the target model to directly transform relevant but incomplete harmful contents into explicitly harmful output with two typical calibration behaviors, as shown in Figure 7 and Figure LABEL:MCC.
(a) Word Calibration
(b) Knowledge Calibration
Figure 7: Two typical behaviors in metaphor calibration: Word Calibration refines terms while preserving structure, while Knowledge Calibration integrates harmful knowledge to enhance professions.
Collaborative Metaphor Calibration.
Metaphor Calibration can be decoupled from jailbreak attacks to crowdsourcing generating, thus isolating high-risk operations for harmful content obtaining stably and discreetly. We analyze the performance of different open-source LLMs in calibrating the benign content from Metaphor-Driven Thinking. Specifically, we keep the target model in MDT as GPT-4o-mini and calibrate its output to toxicity using different LLMs on AdvBench. As the experimental results in Table 3 demonstrated, those open-source LLMs can easily poison benign content into high-quality toxic forms using metaphorical relations, which means that AVATAR can generate harmful content in a distributed manner to evade detection.
Model Calibration Metrics
TP HPR (%) $\Delta$HP CSR (%)
GLM4-9B 11.96$\pm$7.89 71.42 3.24$\pm$1.28 98.00
Qwen2.5-7B 28.15$\pm$11.28 81.23 5.03$\pm$3.52 96.00
Qwen2.5-32B 22.35$\pm$11.62 92.57 9.74$\pm$4.43 96.00
Llama3-8B 14.74$\pm$6.14 77.34 3.67$\pm$0.82 64.00
Table 3: Metaphor calibration performance across open-source LLMs on poisoning GPT-4o-mini’s benign analysis on AdvBench. The following metrics describe the toxicity of calibrated content: 1) TP, Total Phrases, 2) HPR, Harmful Phrase Rate (%), 3) $\Delta$HP, increase in Harmful Phrases, and 4) CSR, Calibration Success Rate.
Defense Assessment.
To further evaluate AVATAR’s ability in manipulating LLMs’ goal prioritization, we use two general adaptive tactics to strengthen LLM’s internal safety, including Adaptive System Prompt (Adapt Sys.) and Targeted Summarization (Tar. Smry.) for adversarial defense. The former reinforces the ethical limits of the given LLM, and the latter uses the given LLM to summarize and rewrite the input prompt for exposing potential harmful intents. As shown in Table 4.2, despite these defense tactics successfully defending against direct jailbreak methods, PAIR, and CoA, they are less effective against AVATAR. Compared with them, AVATAR transforms the malicious intent of jailbreaking as a harmless metaphor analysis task, and strategically couples the harmful content generation with LLM’s reasoning processes. Such a malicious intent decoupling design by reasoning nesting highlights AVATAR’s significant threat to the safety of LLMs. The detailed settings and analysis of external defense are in Appendix.
Defense ASR-GPT (%)
GPT-4o GPT-4o-mini
AVATAR (ours)
No defense 87.20 $\pm$ 8.20 91.33 $\pm$ 3.72
+Adapt Sys. 70.00 $\pm$ 7.48 17.20↓ 85.00 $\pm$ 3.95 6.33↓
+Tar. Smry. 82.00 $\pm$ 8.60 5.20↓ 84.40 $\pm$ 8.29 6.93↓
[…]
No defense 59.20 $\pm$ 7.57 67.60 $\pm$ 5.90
+Adapt Sys. 20.40 $\pm$ 5.18 38.80↓ 28.00 $\pm$ 3.46 39.60↓
+Tar. Smry. 28.80 $\pm$ 6.72 30.40↓ 51.20 $\pm$ 3.63 16.40↓
[…]
No defense 38.80 $\pm$ 7.43 43.60 $\pm$ 5.90
+Adapt Sys. 6.40 $\pm$ 3.85 32.40↓ 10.40 $\pm$ 2.61 33.20↓
+Tar. Smry. 10.00 $\pm$ 2.45 28.80↓ 21.20 $\pm$ 2.28 22.40↓
Table 4: Experimental ASR-GPT (%) of various defense methods on AdvBench, averaging results from 5 repeated experiments.
5 Related Work
Human Value Alignment for LLMs.
Aligning LLMs with human values remains a challenge due to biases in training data and trade-offs between usefulness and safety Zeng et al. (2024); Ding et al. (2024). Approaches such as Reinforcement Learning from Human Feedback (RLHF) Ouyang et al. (2022); Bai et al. (2022) have been proposed to improve fairness Navigli et al. (2023); Gallegos et al. (2024), safety Zou et al. (2023) and eliminate hallucinations Zhang et al. (2023); Lin et al. (2024).
Jailbreak Attacks on LLMs.
Jailbreak attacks threaten the safety alignment mechanisms of LLMs, potentially leading to the generation of harmful content Carlini et al. (2023); Liu et al. (2023); Yi et al. (2024); Li et al. (2025). Our study is inspired by two key methods in black-box attacks: prompt nesting and multi-turn dialogue attacks. 1) Prompt Nesting Attack. Prompt nesting bypasses security features by nesting malicious intents in normal prompts, altering LLMs’ context. DeepInception Li et al. (2023b) exploit nested scenarios, while ReNeLLM Ding et al. (2024) rewrites prompt to jailbreak based on code completion, text continuation, or form-filling tasks. MJP Li et al. (2023a) uses multi-step approaches with contextual contamination to reduce moral constraints, prompting malicious responses. 2) Multi-turn Dialogue Attack. LLMs that are safe in isolated, single-round interactions can be gradually manipulated into generating harmful outputs through multiple rounds of interaction Russinovich et al. (2024); Zhou et al. (2024); Yang et al. (2024); Cheng et al. (2024). Multi-turn dialogue attack leverages the multi-turn nature of conversational interactions to gradually erode an LLM’s content restrictions.
6 Conclusion
In this study, we enhance the understanding of jailbreak attacks by proposing a novel approach, AdVersArial meTAphoR (AVATAR) that manipulates LLMs to generate harmful content through calibrating benign metaphors, rather than generating harmful content from scratch. AVATAR consists of Adversarial Entity Mapping (AEM) and Metaphor-Driven Thinking (MDT). Among them, AEM identifies metaphors by model crowdsourcing, thus enabling the stable acquisition of suitable metaphors before jailbreaking the target LLM. MDT induces the target LLM to analyze the metaphor for harmful content generating, by introducing the metaphor calibration mechanisms, MDT further enhances the efficiency of transforming benign content into harmful forms. Experiments have demonstrated our AVATAR is effective in generating transferable adversarial attacks on multiple advanced LLMs.
Limitations
AVATAR framework generates metaphors using the imagination of LLMs and conducts human-like interaction to prompt LLMs to generate harmful content without complex tricks or sophisticated jailbreak templates. However, the effectiveness of AVATAR is potentially limited by:
Imagination is limited by poor real-world experience.
LLMs lack direct real-world experience and their understanding of harmful entities is predominantly textual. While this vulnerability enables our metaphor-based attack, it also limits the attack’s ability to generate truly realistic and contextually appropriate harmful attacks. To alleviate this issue, we introduce the crowdsourcing strategy to identify the nice metaphors.
Entities Extracting is disturbed by safe alignment.
The safety alignment mechanisms in existing open-source and closed-source LLMs limit their deeper analysis of harmful content for better toxic entity extraction. During the Adversarial Entity Mapping stage, the attack model may extract a mix of harmful and safe entities, e.g., ethical consideration legal consequences, from the harmful query, potentially interfering with subsequent metaphorical content generation. The effectiveness of in-context learning in mitigating this issue is limited. To alleviate this issue, we introduce the word filtering mechanism to remove safe entities during the entity extracting.
Lack of Adaptive Metaphor Switching.
AVATAR currently lacks a mechanism to adaptively switch metaphors for inducing harmful content generation during the metaphor-driven thinking stage. This limitation may reduce the flexibility and effectiveness of the attack in some difficult queries.